# Generative AI Agents for Satellite Networks through a Mixture of Experts Transmission

### v1

@@ III-C4 Generation Strategies
During inference, RAG-Sequence models adopt distinct approximations to $p_{\rm Seq}(d|q)$. This paradigm requires to search on each document $z$, merging hypotheses across documents to form a candidate set. The likelihood $p_{\rm Seq}(d|q)$ is then deduced by aggregating generator probabilities $p_{\theta}(d_{i}|q,z,d_{1:i-1})$ across the document set, thereby generating rational and accurate answers about satellite communications modeling according to user descriptions.
@@ IV Proposed PPO with MoE Approach
As demonstrated in Proposition 1, due to existence of heterogeneous variables and complexity of the optimization environment, Problem (23) is an NP-hard problem, which poses a significant challenge in finding reasonable solutions. Therefore, in this section, we introduce an MoE-PPO approach.
@@ IV-A1 Architecture of the MoE Model
MoE is the integration of multiple specialized sub-networks (i.e., “experts") under the guidance of a central “gating network" [41]. This design consists of multiple expert networks ($E_{1},E_{2},...,E_{I}$), each with unique parameters and running simultaneously. The central gating network labeled $G_{\sigma}$ generates a weight vector that enables the model to direct attention to the most relevant experts based on the current input.
@@ IV-B3 Objective Function with MOE
To align the MOE framework within the PPO, the surrogate objective function is adapted to accommodate the MoE while maintaining the core principles of PPO, which is expressed by
@@ IV-C3 Reward Function
1 $\mathbf{Input:}$ Episodes, exploration steps, learning rate, channel status information, number of experts;
@@ 5             Each expert makes its own action $a^{{}^{\prime}}_{t}$ based on sub-policies;
6             Aggregate actions $a_{t}=\{\textbf{w}^{\rm p}_{k},\textbf{w}^{\rm c},c_{k}\}$ based on the current state $s_{t}$ and gating function;
7             Get the current reward $r_{t}$;
@@ 11             Sample $B$ mini-batch of transitions from $\mathcal{D}$ from all experts;
12             Calculate advantage function $A(s,a)$ for all experts and get the maximization through Back-propagation maximization approach;
13             Update actor network $\theta_{A}$ and critic network $\theta_{C}$ via SGD according to (34) and (35);
@@ V-B1 Satellite Communications Modeling
First, we evaluate the effectiveness of the proposed generative AI agent in customizing satellite communication models. As shown in Fig. 4, the users apply the role assignment strategy at the beginning, enabling the LLM to recall pre-trained satellite knowledge by asking it to act as satellite experts. Then, the users describe the requirements for the satellite communication model from four aspects. We can observe that our agent can precisely capture critical semantic keys from user natural language descriptions and call the corresponding RAG database. Then, the LLM can leverage the retrieved knowledge to generate coherent answers about satellite communications modeling according to user requests. Note that even though the user query is vaGGU, the semantic router can still find the most relevant sub-block by semantic similarity. For instance, the agent precisely routes to the time-varying channel without explicit mention by users. In this case of Fig. 4, the user’s requirement corresponds to a multiple-satellite network using RSMA protocol and time-varying channels, with the goal of maximizing the system EE. We can observe that the AI agent successfully constructs the entire model using 6 rounds of interactions, reaching the minimum for accomplishing five-step modeling (one extra round is for role assignment). Moreover, with the varying network conditions, the user requirement may change over time. Our agent can associate the requirement change to the parts of the models and revise such parts automatically, thus avoiding human-caused errors. As illustrated in Fig. 4, the agent precisely updates the channel model from Rayleigh fading to Rician when the network dynamics decrease.
@@ V-B2 Generation Accuracy
Next, we explore the accuracy of the generated model under different agent settings. Specifically, we adjust two hyperparameters in RAG, namely chunk size and number of chunks [^2] Chunk is the unit of embedding segments. To increase the retrieval efficiency, in each sub-block, the agent splits the expert knowledge embedding into multiple chunks [37].. The former means the number of tokens in each chunk; the latter means the number of chunks that can be retrieved at one time. Note that generation accuracy is defined as $n_{c}/5$, which means the number of correct answers among all five queries in Fig. 4. Fig. 5 demonstrates that low chunk size and number lead to poor generation accuracy. This is because the RAG system cannot fetch expertise to support answer generation. However, if retrieving too much redundant expertise, the LLM of the agent can hardly analyze it effectively, resulting in poor accuracy as well. Therefore, we set chunk size and number to be 500 and 5, respectively, which ensures that our agent can generate precise models with the minimum number of retrieved tokens (2500 in total).
@@ V-C1 The Convergence of PPO with MoE Approach
Fig. 3 illustrates the convergence behavior of the proposed MoE-PPO approach in comparison to some different benchmarks, such as the traditional PPO, greedy, and random selection strategies. The figure reveals that both the MoE-PPO and the standard PPO approaches demonstrate convergence as the number of episodes increases. Notably, MoE-PPO consistently surpasses the performance of PPO, achieving higher sum rates in each iteration. Additionally, the MoE-PPO approach is observed to converge at around 3 million episodes, in contrast to PPO, which requires approximately 5 million episodes. This represents a significant improvement in convergence speed by about 40%. The performance enhancements of MoE-PPO can be attributed to the integration of the MoE framework, which utilizes expert policy specialization and adaptive policy weights to more effectively navigate the policy space. To further validate the effectiveness of MoE-PPO, its comparison with the greedy and random baselines after convergence underscores a significant improvement (i.e., about 210% and 38%). This improvement is facilitated by MoE-PPO’s enhanced collaborative capabilities and more efficient utilization of the state-action space using gating network.
@@ V-C2 Different Number of Experts
V-C3 Adaptability of the MoE-PPO approach
Fig. 8 presents the achievable sum rate obtained by RSMA and SDMA as the number of transmit antennas $N_{t}$. It shows that the achievable sum rate increases with the number of $N_{t}$ for all considered protocols and approachs. This increment is attributable to the additional spatial degrees of freedom provided by more antennas, which help mitigate interference and enhance system capacity. Moreover, it is observed that the system performance under RSMA outperforms that under SDMA (about 42.6% gain). This superiority of RSMA can be ascribed to the employment of successive interference cancellation (SIC) techniques, which effectively remove a part of the interference. Additionally, under each antenna configuration, our MoE-PPO approach always surpasses the PPO approach, which is consistent with the previous simulations.
@@ V-C3 Adaptability of the MoE-PPO approach
It is observed that whether maximizing EE or minimizing power consumpution, our proposed MoE-PPO approach always outperforms the traditional PPO (about 10% and 12%) with the increment of episodes. The reason is that the MoE structure is able to efficiently allocate computing resources because each expert can operate in parallel to solve different variables in which they specialize. This parallelism facilitates deep exploration of the state space, resulting in better policy updates at each stage. Next, it also demonstrates the adaptability of the proposed MoE-PPO approach, confirming its ability to meet different channel fulfillment and optimization goals.

### v2

@@ III-C4 Generation Strategies
During inference, RAG-Sequence models adopt distinct approximations to $p_{\rm Seq}(m|q)$. This paradigm requires searching on each document $z$, merging hypotheses across documents to form a candidate set. The likelihood $p_{\rm Seq}(m|q)$ is then deduced by aggregating generator probabilities $p_{\theta}(m_{i}|q,z,m_{1:i-1})$ across the document set, thereby generating rational and accurate answers about satellite communications modeling according to user descriptions/queries. In summary, Algorithm 1 illustrates the process of performing customized satellite communications modeling via the proposed generative AI agent.
@@ IV Proposed PPO with MoE Approach
As demonstrated in Proposition 1, due to the existence of heterogeneous variables and the complexity of the optimization environment, Problem (23) is an NP-hard problem, which poses a significant challenge in finding reasonable solutions. Therefore, in this section, we introduce an MoE-PPO approach. Note that generative AI agent can generate a variety of optimization problems, we focus on addressing one of the most challenging formulated problems, i.e., (23), constructed by the generative AI agent. For other simpler formulated problems that the proposed AI agent constructs, fewer experts and a more straightforward action space can be used.
@@ IV-A1 Architecture of the MoE Model
MoE is the integration of multiple specialized sub-networks (i.e., “experts”) under the guidance of a central “gating network” [41]. This design consists of multiple expert networks ($E_{1},E_{2},...,E_{I}$), each with unique parameters and running simultaneously. The central gating network labeled $G_{\sigma}$ generates a weight vector that enables the model to direct attention to the most relevant experts based on the current input.
@@ IV-B3 Objective Function with MOE
To align the MOE framework within the PPO, the surrogate objective function is adapted to accommodate the MoE while maintaining the core principles of PPO [42], which is expressed by
@@ IV-B4 Policy and Critic Update Mechanisms
Remark 4.
Note that, in our proposed MoE-PPO approach, we use a back-propagation maximization method to update the experts. This method uses the gradient of the most advantageous function value across all experts to update each expert, effectively allowing the best-performing expert to guide the improvement of the other experts. Additionally, the gating network dynamically adjusts weights, ensuring that the influence of poorly performing experts is mitigated. By doing so, the impact of the worst-performing experts on the whole network is significantly minimized. Such the back-propagation maximization method and gating network have been adopted in existing studies, e.g., [43] and [44].
@@ IV-C3 Reward Function
1 $\mathbf{Input:}$ Episodes, exploration steps, update episodes, learning rate, channel status information, number of experts;
@@ 5             Each expert makes its own action $a^{{}^{\prime}}_{t}$ based on sub-policies;
6             Aggregate actions $a_{t}=\{\textbf{w}^{\rm p}_{k},\textbf{w}^{\rm c},c_{k}\}$ based on the current state $s_{t}$ and gating mechanism;
7             Obtain the current reward $r_{t}$;
@@ 11             Sample $B$ mini-batch of transitions from $\mathcal{D}$ from all experts;
12             Update actor network $\theta_{A}$ and critic network $\theta_{C}$ via SGD according to (34) and (35) for half episodes of total update episodes;
13             Calculate advantage function $A(s,a)$ for all experts and obtain the maximization value $A^{{}^{\prime}}(s,a)$ through Back-propagation maximization approach;
14             Update actor network $\theta_{A}$ and critic network $\theta_{C}$ via SGD and $A^{{}^{\prime}}(s,a)$ according to (34) and (35) for half episodes of total update episodes;
@@ V-B1 Satellite Communications Modeling
First, we evaluate the effectiveness of the proposed generative AI agent in customizing satellite communication models. As shown in Fig. 4, the users apply the role assignment strategy at the beginning, enabling the LLM to recall pre-trained satellite knowledge by asking it to act as satellite experts. Then, the users describe the requirements for the satellite communication model from four aspects. We can observe that our agent can precisely capture critical semantic keys from user natural language descriptions and call the corresponding RAG database. Then, the LLM can leverage the retrieved knowledge to generate coherent answers about satellite communications modeling according to user requests. Note that even though the user query is vague, the semantic router can still find the most relevant sub-block by semantic similarity. For instance, the agent precisely routes to the time-varying channel without explicit mention by users. In this case of Fig. 4, the user’s requirement corresponds to a multiple-satellite network using RSMA protocol and time-varying channels, with the goal of maximizing the system EE. We can observe that the AI agent successfully constructs the entire model using 6 rounds of interactions, reaching the minimum for accomplishing five-step modeling (one extra round is for role assignment). Moreover, with the varying network conditions, the user requirement may change over time. Our agent can associate the requirement change to the parts of the models and revise such parts automatically, thus avoiding human-caused errors. As illustrated in Fig. 4, the agent precisely updates the channel model from Rayleigh fading to Rician when the network dynamics decrease.
@@ V-B2 Generation Accuracy
Next, we evaluate the accuracy of the generated model under different agent settings. For our proposed generative AI agent, it leverages RAG to help human users perform customized satellite communications modeling. Expert modeling knowledge is fed to the agent through chunked embeddings. If the correct knowledge chunk [^2] Chunk is the unit of embedding segments. To increase the retrieval efficiency, in each sub-block, the agent splits the expert knowledge embedding into multiple chunks [14]. that corresponds to the specific user-described scenario is retrieved, the generated answer can be authentic and rational. Therefore, to check the authenticity and rationality of the generated content, the metric named retrieval rate (RR) [49] can be utilized as accuracy. Specifically, RR indicates the ratio that the generative AI agent successfully retrieves the correct knowledge chunk and constructs the answer. Here, we adjust two hyperparameters in RAG, namely the chunk size and the number of chunks. The former means the number of tokens in each chunk; the latter means the number of chunks that can be retrieved at one time. Note that generation accuracy is defined as $n_{c}/5$, which means the number of correct answers among all five queries in Fig. 4. Fig. 5 demonstrates that low chunk size and number lead to poor generation accuracy. This is because the RAG system cannot fetch expertise to support answer generation. However, if retrieving too much redundant expertise, the LLM of the agent can hardly analyze it effectively, resulting in poor accuracy as well. Therefore, we set chunk size and number to be 500 and 5, respectively, which ensures that our agent can generate precise models with the minimum number of retrieved tokens (2500 in total).
@@ V-C The Effectiveness of PPO with MoE Approach
Figure 8: The effectiveness of MoE-PPO under different number of LGUs.
@@ V-C1 The Convergence of PPO with MoE Approach
Fig. 6 illustrates the convergence behavior of the proposed MoE-PPO approach in comparison to some different benchmarks, such as the traditional PPO, greedy, and random selection strategies. The figure reveals that both the MoE-PPO and the standard PPO approaches demonstrate convergence as the number of episodes increases. It shows that MoE-PPO consistently surpasses the performance of PPO (i.e., about 5%), achieving higher sum rates in each iteration. The performance enhancements of MoE-PPO can be attributed to the integration of the MoE framework, which utilizes expert policy specialization and adaptive policy weights to more effectively navigate the policy space. To further validate the effectiveness of MoE-PPO, its comparison with the greedy and random baselines after convergence underscores a significant improvement (i.e., about 220% and 38%). This improvement is facilitated by MoE-PPO’s enhanced collaborative capabilities and more efficient utilization of the state-action space using gating network.
@@ V-C2 Different Number of Experts
V-C3 Different number of LGUs
Fig. 8 illustrates the achievable sum rate as a function of the number of LGUs. It is observed that the achievable sum rate decreases as the number of users increases for both the proposed MoE-PPO approach and the pure PPO approach. This decline occurs because, with increasing users, the approach must allocate more communication resources to users with poor channel quality to meet constraints, reducing the overall sum rate. Moreover, it is evident that the proposed MoE-PPO method consistently outperforms the PPO method, where the MoE-PPO approach achieves an 8.3% higher sum rate compared to the PPO approach when there are 2 LGUs. This superiority is due to the MoE’s ability to leverage multiple specialized experts to handle different aspects of the optimization problem more effectively. Specifically, MoE-PPO can dynamically select the most relevant experts, optimizing resource allocation more efficiently than a single comprehensive network.
V-D The Adaptability of the generative AI agent and MoE-PPO approach
V-D1 MoE-PPO approach for different access protocols from generative AI agent construction
Fig. 9 presents the achievable sum rate obtained by RSMA and SDMA versus the number of transmit antennas $N_{t}$. It shows that the achievable sum rate increases with the number of $N_{t}$ for all considered protocols and approaches. This increment is attributable to the additional spatial degrees of freedom provided by more antennas, which help mitigate interference and enhance system capacity. Moreover, it is observed that the system performance under RSMA outperforms that under SDMA (about 42.6% gain). This superiority of RSMA can be ascribed to the employment of successive interference cancellation (SIC) techniques, which effectively remove a part of the interference. Additionally, under each antenna configuration, our MoE-PPO approach always surpasses the PPO approach, which is consistent with the previous simulations.
@@ V-D1 MoE-PPO approach for different access protocols from generative AI agent construction
V-D2 MoE-PPO approach for different optimization goals from generative AI agent construction
@@
It is observed that whether maximizing EE or minimizing power consumpution, our proposed MoE-PPO approach always outperforms the traditional PPO (about 10% and 12%) with the increment of episodes. The reason is that the MoE structure is able to efficiently allocate computing resources because each expert can operate in parallel to solve different variables in which they specialize. This parallelism facilitates deep exploration of the state space, resulting in better policy updates at each stage. Next, it also demonstrates the adaptability of the proposed MoE-PPO approach, confirming its ability to meet different channel conditions and optimization goals.